# Transverse Instability of Stokes Waves at Finite Depth

### v1

@@ Abstract.
A Stokes wave is a traveling free-surface periodic water wave that is constant in the direction transverse to the direction of propagation. In 1981 McLean discovered via numerical methods that Stokes waves are unstable with respect to transverse perturbations. In [17] for the case of infinite depth we proved rigorously that the spectrum of the water wave system linearized at small Stokes waves, with respect to transverse perturbations, contains unstable eigenvalues lying approximately on an ellipse. In this paper we consider the case of finite depth and prove that the same spectral instability result holds for all but finitely many values of the depth. The computations and some aspects of the theory are considerably more complicated in the finite depth case.
@@ 1. Introduction
We consider classical water waves that are irrotational, incompressible, and inviscid. The water lies below an unknown free surface $S$. Such waves have been studied for over two centuries, notably by Stokes [36]. A Stokes wave is a two-dimensional steady wave traveling in a fixed horizontal direction at a fixed speed $c$. It has been known for a century that a curve of small-amplitude Stokes waves exists [31, 27, 37]. Several decades ago it was proven that the curve extends to large amplitudes as well [24].
@@
For background and references on longitudinal (modulational and high frequency) instabilities, we refer to [17]. To bring the references up to date, we mention some very recent work on longitudinal instability by Berti et al [8] and on near-extreme Stokes waves by Deconinck et al [20]. It is important to note that there are several other models of water waves for which the transverse instability has been studied rigorously, for instance for gravity-capillary waves.
As in [17] we now specify the parameters of our problem. Let $x$ and $y$ denote the horizontal variables and $z$ the vertical one. The problem depends in a significant way on the depth $h$. Consider the curve of Stokes waves traveling in the $x$-direction and with a given period, say $2\pi$ without loss of generality. This curve is parametrized by a small parameter $\varepsilon$ which represents the wave amplitude of the Stokes waves. Such a steady wave can be described in the moving $(x,z)$ plane (where $x-ct$ is replaced by $x$) by its free surface $S=\{(x,y,z)\ |\ z=\eta^{*}(x;\varepsilon)\}$ and by its velocity potential $\psi^{*}(x;\varepsilon)$ restricted to $S$.
We take the perturbation of $\eta^{*}$ to have the form $\overline{\eta}(x)e^{\lambda t+i\alpha y}$, where $\overline{\eta}$ has the same period $2\pi$ as the Stokes wave, $\lambda\in\mathbb{C}$ is the growth rate of the perturbation, and $\alpha\in\mathbb{R}$ is the transverse wave number of the perturbation. The problem is to find at least one value of $\alpha$ that leads to instability, that is, $\text{Re}\lambda>0$. After linearizing the nonlinear water wave system about a Stokes wave, introducing a “good-unknown,” and performing a conformal mapping change of variables, we find that the exponents $\lambda$ are eigenvalues of a linear Hamiltonian operator $\mathcal{L}_{h,\varepsilon,\beta}$, where $\beta=\alpha^{2}$. Of course, the operator depends in a significant way on the depth $h$.
@@
It is clear that the limit as $h\to\infty$ is simpler. Motivated by [30], we determine a resonant transverse wave number $\alpha_{*}$ so that the unperturbed operator $\mathcal{L}_{h,0,\beta_{*}}$ with $\beta_{*}=\alpha_{*}^{2}$ has an imaginary double eigenvalue $i\sigma$. This eigenvalue corresponds to the lowest possible resonance that generates a Type II transverse instability according to McLean [30], of which there are infinitely many higher-order resonances that potentially generate higher-order transverse instabilities. We expect however that higher-order transverse instabilities have smaller growth rates for small Stokes waves.
@@
In Section 2 we introduce the Stokes waves and proceed with the linearization and the flattening by means of a conformal mapping from a strip $\Omega_{\varepsilon}=\mathbb{R}\times(-h_{\varepsilon},0)$ to the two-dimensional fluid domain of the Stokes wave. The main result that is required here is Theorem 2.6, which is devoted to the three-dimensional Dirichlet-Neumann operator $\mathcal{G}_{\varepsilon,\beta}$ that occurs in the linearized operator subject to the two-dimensional conformal mapping, and the proof of its analyticity in $\varepsilon$ and $\delta$. The lack of a trace theorem from $H^{1/2}$ of a domain to its boundary necessitates a careful alternative analysis. Compared to the infinite depth case [17], the proof of analyticity is considerably more difficult due to the dependence on $\varepsilon$ of the strip $\Omega_{\varepsilon}$.
Theorem 1.1 is then reduced to studying the eigenvalues of the linearized operator $\mathcal{L}_{h,\varepsilon,\beta}$, which has a Hamiltonian form and is reversible. Section 3 is devoted to a discussion of the first resonance on the imaginary axis when $\varepsilon=0$, the introduction of Kato’s perturbed basis, and the reduction to the study of the eigenvalues of a reduced $2\times 2$ matrix $\textrm{L}_{h,\varepsilon,\delta}$. In Section 4 we perform extremely tedious expansions of $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ out to third order in both $\varepsilon$ and $\delta$, which are considerably more tedious than the corresponding expansions in [17, 32, 4]. We use the expansions of $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ to compute the expansions of the Kato basis vectors and of the matrix $\textrm{L}_{h,\varepsilon,\delta}$. Finally in Section 5 we analyze the leading terms in the characteristic discriminant of $\textrm{L}_{h,\varepsilon,\delta}$. An important step is to prove that a key coefficient, which we call $b_{3,0}$, does not vanish. It is here that we must exclude the exceptional depths. By determining the precise asymptotics of $b_{3,0}$ as $h\to 0^{+}$ and $h\to\infty$, we are able to prove that there are at least one but at most only finitely many such exceptional depths. This concludes the proof of Theorem 1.1. The full equation that defines the isola of eigenvalues up to $O(\varepsilon^{4})$ is given in Corollary 5.4. Finally, we prove a trace result in Appendix A.1, which is used in the proof of the analyticity of the operator $\mathcal{G}_{\varepsilon,\beta}$.
@@ 2.1. Stokes waves
where $P$ is the Bernoulli constant and $g>0$ is the constant acceleration due to gravity. The second equation is Bernoulli’s, which follows from the pressure being constant along the free surface; the third equation expresses the kinematic boundary condition that particles on the surface remain there; the last condition asserts that there is no flux of water out of the bottom boundary $\{z=-h\}$. For convenience we will take $P=0$.
@@
By the change of variables
$c\to\sqrt{g}{c},\quad\eta(x,y,t)\to\eta(x,y,\sqrt{g}t),\quad\psi(x,y,t)\to\sqrt{g}\psi(,x,y,\sqrt{g}t),$
we henceforth assume without loss of generality that $g=1$.
By a Stokes wave we mean a periodic steady solution of (2.5) that is independent of $y$. Without loss of generality, we henceforth consider $2\pi$-periodic Stokes waves. As has been known for over a century, there exists a curve of small Stokes waves parametrized analytically by the small amplitude $\varepsilon$.
@@
  • (ii) $\eta^{*}(\cdot;\varepsilon)$ is even and has average zero, and $\psi^{*}(\cdot;\varepsilon)$ is odd.
@@ 2.2. Transverse perturbations, linearization, and transformations
$B(\eta)\psi:=\frac{G(\eta)\psi+\nabla\eta\cdot\nabla\psi}{1+|\nabla\eta|^{2}},\quad V(\eta)\psi:=\nabla\psi-B\nabla\psi.$ (2.15)
@@
1. Assume first that there exist $h_{\varepsilon}>0$ and a conformal mapping $(X,Z):\Omega\to\Omega_{\varepsilon}$ such that (i) and (ii) hold. We will prove (iii) and (iv). For brevity, denote $f(x)=X(x,0)-x$. We have $\Delta X=0$ and $\partial_{z}X(x,-h_{\varepsilon})=0$ as well as $\Delta Z=0$ and $Z(x,-h_{\varepsilon})=-h$. By solving the Laplace problem for $X$ and using the Cauchy-Riemann equations, we find
@@
Substituting this formula for $\hat{f}$ into (2.28) and (2.29), and noting that $X(x,z)-x$ is odd in $x$, we obtain (2.25) and (2.26).
@@ 2.3. The transformed Dirichlet-Neumann operator
Remark 2.7.
In the infinite depth case, Theorem 2.6 was proven in Theorem 2.7, [17]. The assertions (i) and (iii) can be obtained similarly to the infinite depth case. On the other hand, the analyticity of $\mathcal{G}_{\varepsilon,\beta}$ in $(\varepsilon,\beta)$ is more difficult to establish due to the dependence on $\varepsilon$ of the flat domain $\Omega_{\varepsilon}$. We will be led to proving the analyticity of solutions to the second-order elliptic problem (2.44) where the highest order term has an $\varepsilon$-dependent coefficient. Moreover, the case $s\in[-\frac{1}{2},1]$ will be deduced by interpolation from the cases $s=-\frac{1}{2}$ and $s>1$, where $s=-\frac{1}{2}$ requires the nonstandard trace result in Lemma A.2.
Proof.
We assume without loss of generality that $\alpha>0$ and denote $\mathbb{T}^{2}_{\alpha}=\mathbb{T}\times(\mathbb{R}/2\pi\alpha\mathbb{Z})$. Let $(X,Z):\Omega_{\varepsilon}\to\Omega$ be the Riemann mapping in Theorem 2.4. It has the Jacobian $\mathcal{J}=\mathcal{J}(x,z,\varepsilon)=|\partial_{x}X|^{2}+|\partial_{z}X|^{2}$. For an arbitrary function $f\in H^{\frac{1}{2}}(\mathbb{T})$, we will define the function $\Theta_{f}$ by solving
@@
where $\partial_{z}\Theta_{f}(\cdot,0)$ is the trace $\gamma(\nabla\Theta_{f})$ given in Proposition A.1. As shown in the proof of Theorem 2.7 in [17], for any $s\geq\frac{1}{2}$ there exists $\varepsilon_{0}(s)>0$ such that
@@
The remainder of this proof is devoted to the analyticity (2.35). We fix $s\geq\frac{1}{2}$ unless stated otherwise. We first note that since $\eta^{*}$ is analytic in $\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))$ with values in Sobolev spaces, so are $\zeta$, $(X,Z)-\operatorname{I}$, $\mathcal{J}$, and $h_{\varepsilon}$. For fixed $f\in H^{s}(\mathbb{T})$, we let $\Theta(\varepsilon,\beta)\equiv\Theta(x,z;\varepsilon,\beta)$ be the solution of (2.37), which is posed on $\Omega_{\varepsilon}=\{(x,z):x\in\mathbb{T},~{}z\in(-h_{\varepsilon},0)\}$. Compared to the infinite depth case considered in [17], the extra complication comes from the $\varepsilon$-dependence of the domain $\Omega_{\varepsilon}$. To cope with this we make the simple rescaling
@@
$\begin{cases}\left(\partial_{x}^{2}+\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}^{2}\right)\Psi-\beta\tilde{J}\Psi=0\quad\text{ in }\Omega_{0}\ ,\\ \Psi(x,0)=f(x),\quad\partial_{z}\Psi(x,-h)=0,\end{cases}$ (2.44)
where $\Omega_{0}:=\{(x,z):x\in\mathbb{T},~{}z\in(-h,0)\}$. Thus the $\varepsilon$-dependence of $\Omega_{\varepsilon}$ is now encoded in the coefficient $\frac{h}{h_{\varepsilon}}$ of the highest order term $\partial_{z}^{2}\Psi$. Recalling (2.38) and using the chain rule, we obtain
$\mathcal{G}_{\varepsilon,\beta}f=\frac{h}{h_{\varepsilon}}\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}=\frac{h}{h_{\varepsilon}}\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta)),$ (2.45)
where $\tilde{\gamma}$ is defined in Lemma A.2. We divide the rest of the proof into several steps.
1. We claim that the mapping
$(-\varepsilon_{0}(s),\varepsilon_{0}(s))\times(0,\infty)\ni(\varepsilon,\beta)\mapsto\Psi(\varepsilon,\beta)\in H^{s+\frac{1}{2}}(\Omega_{0})$ (2.46)
is $C^{\infty}$ for $s\geq\frac{1}{2}$.
For sufficiently small $\varepsilon_{0}(s)$, we have $h_{\varepsilon}\geq\frac{1}{2}h$ and $\tilde{J}\geq\frac{1}{2}$. Then, elliptic estimates yield the uniform (in $\varepsilon$) bound
$\|\Psi(\varepsilon,\beta)\|_{H^{s+\frac{1}{2}}(\Omega_{0})}\leq C_{1}(\beta,s)\|f\|_{H^{s}(\mathbb{T})}\quad\forall\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s)),$ (2.47)
where $C_{1}(\beta,s)$ is bounded for $\beta$ in compact subsets of $(0,\infty)$. For fixed $\varepsilon\in(-\varepsilon_{0}(s),\varepsilon_{0}(s))$, the differentiability of $\Psi(\varepsilon,\beta)$ in $\beta$ can be proven analogously to the proof of Theorem 2.7, [17]. On the other hand, for fixed $\beta>0$, the differentiability of $\Psi(\varepsilon,\beta)$ in $\varepsilon$ requires more attention since $\varepsilon$ appears in the highest order term $\frac{h}{h_{\varepsilon}}\partial_{z}^{2}\Psi$. We let $u_{\varepsilon,\beta}$ be the solution of the inhomogeneous problem

### v2

@@ Abstract.
A Stokes wave is a traveling free-surface periodic water wave that is constant in the direction transverse to the direction of propagation. In 1981 McLean discovered via numerical methods that Stokes waves are unstable with respect to transverse perturbations. In [18] for the case of infinite depth we proved rigorously that the spectrum of the water wave system linearized at small Stokes waves, with respect to transverse perturbations, contains unstable eigenvalues lying approximately on an ellipse. In this paper we consider the case of finite depth and prove that the same spectral instability result holds for all but finitely many values of the depth. The computations are considerably more complicated in the finite depth case.
@@ 1. Introduction
We consider classical water waves that are irrotational, incompressible, and inviscid. The water lies below an unknown free surface $S$. Such waves have been studied for over two centuries, notably by Stokes [41]. A Stokes wave is a two-dimensional steady wave traveling in a fixed horizontal direction at a fixed speed $c$. It has been known for a century that a curve of small-amplitude Stokes waves exists [35, 31, 42]. Several decades ago it was proven that the curve extends to large amplitudes as well [28, 27, 43, 3, 39].
@@
For background and references on longitudinal (modulational and high frequency) instabilities, we refer to [18]. To bring the references up to date, we mention some very recent work on longitudinal instability by Berti et al [9] and on near-extreme Stokes waves by Deconinck et al [21]. It is important to note that there are several other models of water waves for which the transverse instability has been studied rigorously, for instance for gravity-capillary waves [40, 24].
As in [18] we now specify the parameters of our problem. Let $x$ and $y$ denote the horizontal variables and $z$ the vertical one. The problem depends in a significant way on the depth $h$. Consider the curve of Stokes waves traveling in the $x$-direction and with a given period, say $2\pi$ without loss of generality. This curve is parametrized by a small parameter $\varepsilon$ which represents the wave amplitude of the Stokes waves. Such a steady wave can be described in the moving $(x,z)$ plane (where xctx-ct is replaced by $x$) by its free surface $S=\{(x,y,z)\ |\ z=\eta^{*}(x;\varepsilon)\}$ and by its velocity potential $\psi^{*}(x;\varepsilon)$ restricted to $S$.
We take the perturbation of $\eta^{*}$ to have the form $\overline{\eta}(x)e^{\lambda t+i\alpha y}$, where $\overline{\eta}$ has the same period $2\pi$ as the Stokes wave, $\lambda\in\mathbb{C}$ is the growth rate of the perturbation, and $\alpha\in\mathbb{R}$ is the transverse wave number of the perturbation. In other words, we consider purely transverse perturbations, leaving out any effect of modulational instabilities. The problem is to find at least one value of $\alpha$ that leads to instability, that is, $\text{Re}\lambda>0$. After linearizing the nonlinear water wave system about a Stokes wave, introducing a “good-unknown,” and performing a conformal mapping change of variables, we find that the exponents $\lambda$ are eigenvalues of a linear Hamiltonian operator $\mathcal{L}_{h,\varepsilon,\beta}$, where $\beta=\alpha^{2}$. Of course, the operator depends in a significant way on the depth $h$.
@@
It is clear that the limit as $h\to\infty$ is simpler. Motivated by [34], we determine a resonant transverse wave number $\alpha_{*}$ so that the unperturbed operator $\mathcal{L}_{h,0,\beta_{*}}$ with $\beta_{*}=\alpha_{*}^{2}$ has an imaginary double eigenvalue $i\sigma$. This eigenvalue corresponds to the lowest possible resonance ($m=1$ in (3.4)) that generates a Type II transverse instability according to McLean [34], of which there are infinitely many higher-order resonances that potentially generate higher-order transverse instabilities. We expect however that higher-order transverse instabilities have smaller growth rates for small Stokes waves. More precisely, we conjecture that the $m^{th}$ resonance in (3.4) would lead to instabilities at order $\varepsilon^{2m+1}$.
@@
Shortly after the appearance of the preprint version of this paper, the preprint [23] of Jiao-Rodrigues-Yang proved the oblique instability of small Stokes waves when additional localized perturbations in the longitudinal direction are included. They thereby obtained an $\varepsilon^{2}$-order instability, which is expected from the dominant modulational instability [36, 5].
Remark 1.4.
@@
In Section 2 we introduce the Stokes waves and proceed with the linearization and the flattening by means of a conformal mapping from a strip $\Omega_{\varepsilon}=\mathbb{R}\times(-h_{\varepsilon},0)$ to the two-dimensional fluid domain of the Stokes wave. The main result that is required here is Theorem 2.6, which is devoted to the three-dimensional Dirichlet-Neumann operator $\mathcal{G}_{\varepsilon,\beta}$ that occurs in the linearized operator subject to the two-dimensional conformal mapping, and the proof of its analyticity in $\varepsilon$ and $\delta$. The analyticity is proven by an approach inspired by Groves [22] and is significantly shorter than our direct approach in [18].
Theorem 1.1 is then reduced to studying the eigenvalues of the linearized operator $\mathcal{L}_{h,\varepsilon,\beta}$, which has a Hamiltonian form and is reversible. Section 3 is devoted to a discussion of the first resonance on the imaginary axis when $\varepsilon=0$, the introduction of Kato’s perturbed basis, and the reduction to the study of the eigenvalues of a reduced $2\times 2$ matrix $\textrm{L}_{h,\varepsilon,\delta}$. In Section 4 we perform expansions of $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ out to third order in both $\varepsilon$ and $\delta$, which are considerably more tedious than the corresponding expansions in [18, 36, 5]. We use the expansions of $\mathcal{L}_{h,\varepsilon,\beta_{*}+\delta}$ to compute the expansions of the Kato basis vectors and of the matrix $\textrm{L}_{h,\varepsilon,\delta}$. Finally in Section 5 we analyze the leading terms in the characteristic discriminant of $\textrm{L}_{h,\varepsilon,\delta}$. An important step is to prove that a key coefficient, which we call $b_{3,0}$, does not vanish. It is here that we must exclude the exceptional depths. By determining the precise asymptotics of $b_{3,0}$ as $h\to 0^{+}$ and $h\to\infty$, we are able to prove that there are at least one but at most only finitely many such exceptional depths. This concludes the proof of Theorem 1.1. The full equation that defines the isola of eigenvalues up to $O(\varepsilon^{4})$ is given in Corollary 5.4.
@@ 2.1. Stokes waves
where $P$ is the Bernoulli constant and $g>0$ is the constant acceleration due to gravity. The second equation is Bernoulli’s, which follows from the pressure being constant along the free surface; the third equation expresses the kinematic boundary condition that particles on the surface remain there; the last condition asserts that there is no flux of water out of the bottom boundary $\{z=-h\}$. Without loss of generality we will take $P=0$.
@@
For any $a>0$, if we make the change of variables
$\eta(x,y,t)=\frac{1}{a}\tilde{\eta}(ax,ay,\sqrt{g}t),\quad\psi(x,y)=\frac{\sqrt{g}}{a\sqrt{a}}\tilde{\psi}(ax,ay,\sqrt{g}t),$
[…]
By a Stokes wave we mean a periodic steady solution of (2.5) that is independent of $y$. In view of the aforementioned scaling, without loss of generality, we henceforth consider $2\pi$-periodic Stokes waves. As has been known for over a century, there exists a curve of small Stokes waves parametrized analytically by the small amplitude $\varepsilon$.
@@
  • (ii) $\eta^{*}(\cdot;\varepsilon)$ is even, and $\psi^{*}(\cdot;\varepsilon)$ is odd.
@@ 2.2. Transverse perturbations, linearization, and transformations
$B(\eta)\psi:=\frac{G(\eta)\psi+\nabla\eta\cdot\nabla\psi}{1+|\nabla\eta|^{2}},\quad V(\eta)\psi:=\nabla\psi-\nabla\eta B(\eta)\psi.$ (2.15)
@@
1. Assume first that there exist $h_{\varepsilon}>0$ and a conformal mapping $(X,Z):\Omega\to\Omega_{\varepsilon}$ such that (i) and (ii) hold. We will prove (iii) and (iv). For brevity, denote f(x)=X(x,0)xf(x)=X(x,0)-x. We have $\Delta X=0$ and $\partial_{z}X(x,-h_{\varepsilon})=0$ as well as $\Delta Z=0$ and $Z(x,-h_{\varepsilon})=-h$. By solving the Laplace problem for $X$ and using the Cauchy-Riemann equations, we find
@@
Substituting this formula for $\hat{f}$ into (2.28) and (2.29), and noting that X(x,z)xX(x,z)-x is odd in $x$, we obtain (2.25) and (2.26).
@@ 2.3. The transformed Dirichlet-Neumann operator
In the infinite depth case, Theorem 2.6 was proven in Theorem 2.7, [18]. The assertions (i) and (iii) can be obtained similarly to the infinite depth case. Regarding the analyticity of $G_{\varepsilon,\beta}$ in $(\varepsilon,\beta)$, the direct proof in [18] can be adapted to the finite depth case. However, the low regularity case $s=\frac{1}{2}$ would require a careful treatment, due to the $\varepsilon$-dependence of the conformal domain $\{-h_{\varepsilon}<z<0\}$ and the trace operator for vector fields not being in $H_{\operatorname{div}}$ (see (2.37)) which arise if one changes $z\mapsto z\frac{h}{h_{\varepsilon}}$ to cope with $h_{\varepsilon}$. We will provide an alternative, shorter proof using the analytic implicit function theorem, in the spirit of the proof of Lemma 2.2 in [22]. We will need the following classical trace operator.
Proposition 2.7 (Section 3.2, Chapter IV, [10]).
$H_{\operatorname{div}}(U)=\{u\in L^{2}(U;\mathbb{R}^{n}):\operatorname{div}u\in L^{2}(U)\}.$ (2.37)
There exists a bounded linear operator $\gamma:H_{\operatorname{div}}(U)\to H^{-\frac{1}{2}}(\partial U)$ such that $\gamma(u)=u\cdot\nu$ for $u\in C^{\infty}_{0}(\overline{U})$, where $\nu$ is the unit outward normal to $\partial U$. In particular, there exists a constant $C>0$ such that
$\|\gamma(u)\|_{H^{-\frac{1}{2}}(\partial U)}\leq C\Big(\|u\|_{L^{2}(U)}+\|\operatorname{div}u\|_{L^{2}(U)}\Big)\quad\forall u\in H_{\operatorname{div}}(U).$ (2.38)
$\langle\gamma(u),\gamma_{0}(w)\rangle_{H^{-\frac{1}{2}}(\partial U),H^{\frac{1}{2}}(\partial U)}=\int_{U}u\cdot\nabla w+\int_{U}w\operatorname{div}u$ (2.39)
holds, where $\gamma_{0}(w)$ is the trace of $w$.
Proof of Theorem 2.6
We assume without loss of generality that $\alpha>0$ and denote $\mathbb{T}^{2}_{\alpha}=\mathbb{T}\times(\mathbb{R}/2\pi\alpha\mathbb{Z})$. Let $(X,Z):\Omega_{\varepsilon}\to\Omega$ be the Riemann mapping in Theorem 2.4. It has the Jacobian $\mathcal{J}=\mathcal{J}(x,z,\varepsilon)=|\partial_{x}X|^{2}+|\partial_{z}X|^{2}$, which is analytic in $\varepsilon$ near 0. For an arbitrary function $f\in H^{\frac{1}{2}}(\mathbb{T})$, we will define the function $\Theta_{f}$ by solving
@@
where $\partial_{z}\Theta_{f}(\cdot,0)$ is the trace $\gamma(\nabla\Theta_{f})$ given in Proposition 2.7. As shown in the proof of Theorem 2.7 in [18], for any $s\geq\frac{1}{2}$ there exists $\varepsilon_{0}(s)>0$ such that
@@
The remainder of this proof is devoted to the analyticity (2.35) of $\mathcal{G}_{\varepsilon,\beta}$. We denote
$\Omega_{0}:=\{(x,z):x\in\mathbb{T},~z\in(-h,0)\}.$
We fix $\beta_{0}>0$ and $f\in H^{s}(\mathbb{T})$ with $s\geq\frac{1}{2}$. We first prove that for any $F_{0}\in H^{s-\frac{1}{2}}(\Omega_{0})$ and $G\in(H^{s-\frac{1}{2}}(\Omega_{0}))^{2}$, the problem
$\begin{cases}\Delta_{x,z}v-\beta_{0}v=F_{0}+\operatorname{div}G\quad\text{ in }\Omega_{0},\\ v(x,0)=f(x),\\ \partial_{z}v(x,-h)=G_{2}(x,-h)\end{cases}$ (2.47)
has a unique solution $v\in H^{s+\frac{1}{2}}(\Omega_{0})$ in the following sense: $v|_{\{z=0\}}=f$ in the trace sense and
$\int_{\Omega_{0}}\nabla_{x,z}v\cdot\nabla_{x,z}\varphi+\beta_{0}v\varphi\,dxdz=\int_{\Omega_{0}}-F_{0}\varphi+G\cdot\nabla_{x,z}\varphi\,dxdz\quad\forall\varphi\in H^{1,0}(\Omega_{0}),$ (2.48)
where $H^{1,0}(\Omega_{0})=\{u\in H^{1}(\Omega_{0}):~u|_{\{z=0\}}=0\}$. Note that (2.48) implies $\Delta_{x,z}v-\operatorname{div}G=\operatorname{div}(\nabla_{x,z}v-G)=\beta_{0}v+F_{0}\in L^{2}(\Omega_{0})$ in the sense of distributions, and hence
$\int_{\Omega_{0}}(\nabla_{x,z}v-G)\cdot\nabla_{x,z}\varphi+\operatorname{div}(\nabla_{x,z}v-G)\varphi\,dxdz=0\quad\forall\varphi\in H^{1,0}(\Omega_{0}).$
In view of the Stokes formula (2.39), we deduce that the normal trace of $\nabla_{x,z}v-G$ at $\{z=-h\}$ vanishes in $H^{-\frac{1}{2}}(\mathbb{T})$, justifying the Neumann condition $\partial_{z}v(x,-h)=G_{2}(x,-h)$.
For $s=\frac{1}{2}$, the existence and uniqueness of $v\in H^{1}(\Omega_{0})$ follow from the Lax-Milgram theorem and the lifting operator from $H^{s}(\mathbb{T}^{d})$ to $H^{s+\frac{1}{2}}(\Omega_{0})$. Moreover, we have
$\|v\|_{H^{1}(\Omega_{0})}\leq C(\|F_{0}\|_{L^{2}(\Omega_{0})}+\|G\|_{(L^{2}(\Omega_{0}))^{2}}+\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}).$
For $s>\frac{1}{2}$, it remains to obtain the regularity $v\in H^{s+\frac{1}{2}}(\Omega_{0})$. We decompose $v=v_{1}+v_{2}$, where $v_{1}\in H^{1}(\Omega_{0})$ solves
$\begin{cases}\Delta_{x,z}v_{1}-\beta_{0}v_{1}=F_{0}\quad\text{ in }\Omega_{0},\\ v_{1}(x,0)=f(x),\\ \partial_{z}v_{1}(x,-h)=0,\end{cases}$ (2.49)
and $v_{2}\in H^{1}(\Omega_{0})$ solves
$\begin{cases}\Delta_{x,z}v_{2}-\beta_{0}v_{2}=\operatorname{div}G\quad\text{ in }\Omega_{0},\\ v_{2}(x,0)=0,\\ \partial_{z}v_{2}(x,-h)=G_{2}(x,-h).\end{cases}$ (2.50)
Since $F_{0}\in H^{s+\frac{1}{2}}(\Omega_{0})$ has nonnegative regularity, combining standard elliptic regularity with linear interpolation yields $v_{1}\in H^{s+\frac{1}{2}}(\Omega_{0})$. This is also the case for $v_{2}$ provided $s\geq\frac{3}{2}$, so that $\operatorname{div}G\in H^{s-\frac{3}{2}}(\Omega_{0})$ has nonnegative regularity. Since the mapping $H^{s-\frac{1}{2}}(\Omega_{0})\ni G\mapsto v_{2}\in H^{s+\frac{1}{2}}(\Omega_{0})$ is linear, the remaining case $s\in(\frac{1}{2},\frac{3}{2})$ follows by linear interpolation between $s=\frac{1}{2}$ and $s=\frac{3}{2}$.
We denote the mapping $(F_{0},G,f)\mapsto v$ by
$v=\mathcal{S}(F_{0},G,f).$ (2.51)
Since
$\mathcal{S}(F_{0},G,f)=\mathcal{S}(0,0,f)+\mathcal{S}(F_{0},0,0)+\mathcal{S}(0,G,0)$
and each operator on the right-hand side is linear in its respective argument, it follows that
$\mathcal{S}:H^{s-\frac{1}{2}}(\Omega_{0})\times(H^{s-\frac{1}{2}}(\Omega_{0}))^{2}\times H^{s}(\mathbb{T}^{d})\to H^{s+\frac{1}{2}}(\Omega_{0})\quad\text{is~ analytic}.$ (2.52)
Returning to the problem (2.40), in order to cope with the $\varepsilon$-dependence of $\Omega_{\varepsilon}$, we make the simple rescaling
@@
$\begin{cases}\left(\partial_{x}^{2}+\big(\frac{h}{h_{\varepsilon}}\big)^{2}\partial_{z}^{2}\right)\Psi-\beta\tilde{J}\Psi=0\quad\text{ in }\Omega_{0}\ ,\\ \Psi(x,0)=f(x),\quad\partial_{z}\Psi(x,-h)=0.\end{cases}$ (2.53)
Use of the chain rule in the definition (2.41) of $\mathcal{G}_{\varepsilon,\beta}$ yields
$\tfrac{h_{\varepsilon}}{h}\mathcal{G}_{\varepsilon,\beta}f=\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}.$ (2.54)
We rewrite (2.53) as